The pumunary artery which connects the heart to the lungs has an inner radius of 2.6mm and is 8.5cm long. If the pressure drop between the heart and the lungs is 3*10³N/m², what is the average speed of blood in the pumunary artery? Determine volume flow rate of blood. Assume body temperature is 37°C and the coefficient of viscosity of blood is 4*10³ at 37°C

1
Expert's answer
2022-03-04T08:31:32-0500

"v=\frac{Fl}{\eta S},"

"F=pA=\pi pr^2,"

"v=\frac{\pi pr^2l}{\eta S}\approx 1~\frac ms."